H H-C-HQ) + 20=0(0) O=C=0(a) + 2H-O-Hq) Bond enthalpies: C-H = +424 kJ/mol; C==0 = +805 kJ/mol; 0-H = +464 kJ/mol
Answer #1

Chytron, COLTZH-0-6 AH = suning Clity 4C-HA UX424=+ 1696 kJ/md co2ac07 2x 80s=+1610 kJ I onda eto: 20.6 - 2x 460= + 930 lglim